# Variational Bayesian Bow tie Neural Networks with Shrinkage

### v2

@@ 2.3 Polya-Gamma Data Augmentation
As in Smith et al. (2021), we employ Polya-Gamma data augmentation Polson et al. (2013) to render the model conditionally linear and Gaussian. First, recall the definition of the Polya-Gamma distribution with parameters $b>0$ and $c\in\mathbb{R}$, denoted $\text{PG}(b,c)$. The random variable $X\sim\text{PG}(b,c)$ if
@@
where $\kappa=a-b/2$ and $p(\omega)=PG(\omega|b,0)$. The integral is a Gaussian kernel, thus if $z=\mathbf{w}^{T}\mathbf{x}$, conditioned on the latent variable $\omega$, $\mathbf{w}$ has a Gaussian distribution and conditioned on $\mathbf{w}$, $\omega$ has a PG distribution. While to sample from the PG distribution, one can use the alternating series method of Devroye (2006), all finite moments of the PG random variables are available in closed form and that becomes useful for expectation-maximization or variational Bayes algorithms. Specifically, for $c>0$
@@ 2.4 Augmented Model
The model described in Section 2.1 augmented with stochastic activations $\mathbf{a}=(\mathbf{a}_{n,l})$ and binary activations $\bm{\gamma}=(\bm{\gamma}_{n,l})$ is:
@@
Figure 3: Directed Acyclic Graph of the model.
To ease computations, the covariance matrices are assumed to be diagonal $\bm{\Sigma}_{l}=\text{diag}(\eta_{l,1}^{2},\ldots\eta_{l,D_{l}}^{2})$, and the variances are denoted by $\bm{\eta}_{l}=(\eta_{l,1}^{2},\ldots\eta_{l,D_{l}}^{2})$. Additionally, we assume conjugate priors for the variances $\eta_{l,d}^{2}\overset{iid}{\sim}\text{IG}(\alpha^{h}_{0},\beta^{h}_{0})$ for $l=1,\ldots,L$ and $\eta_{L+1,d}^{2}\overset{iid}{\sim}\text{IG}(\alpha_{0},\beta_{0})$ and for the biases $b_{l,d}\overset{iid}{\sim}\text{N}(0,s^{2}_{0})$. Here, we consider different prior parameters $\alpha_{0}^{h},\beta_{0}^{h}$ for the variance terms associated to the hidden layers in comparison to the prior parameters $\alpha_{0},\beta_{0}$ for the final layer. In particular, $\alpha_{0},\beta_{0}$ are chosen to reflect uncertainty in noise, while $\alpha_{0}^{h},\beta_{0}^{h}$ are chosen so that prior concentrates on small values and realizations of the stochastic activation function are more similar to the ReLU.
@@
×$\times\prod_{n=1}^{N}\prod_{l=1}^{L}\prod_{d=1}^{D_{l}}\exp\left(\frac{\kappa_{n,l,d}z_{n,l,d}}{T}\right)\exp\left(-\frac{\omega_{n,l,d}z_{n,l,d}^{2}}{2T^{2}}\right)p(\omega_{n,l,d})$
×$\times\prod_{n=1}^{N}\prod_{l=1}^{L}\prod_{d=1}^{D_{l}}\text{Bern}\left(\gamma_{n,l,d}\mid\sigma\left(\frac{z_{n,l,d}}{T}\right)\right)$
×$\times\prod_{d=1}^{D_{L+1}}\text{IG}(\eta^{2}_{L+1,d}\mid\alpha_{0},\beta_{0})\times\prod_{l=1}^{L}\prod_{d=1}^{D_{l}}\text{IG}(\eta^{2}_{l,d}\mid\alpha_{0}^{h},\beta_{0}^{h})$
@@ 3.1 Variational Bayes
While Markov chain Monte Carlo is considered the gold-standard tool for approximating posterior distributions in Bayesian modelling due its asymptotic guarantees, MCMC algorithms can be prohibitively slow when the model dimension and sample size are large. An alternative fast approximate Bayesian inference method known as variational inference has gained popularity in the literature Ormerod and Wand (2010); Zhang et al. (2018), due to both the explosion in the amount of data collected and use of highly parametrized models for increased flexibility. VI has been shown to yield reasonably accurate approximations in several problems as well as desirable frequentist properties. Namely, consistency and asymptotic normality of VI are studied in Wang and Blei (2019), theoretical guarantees for optimal contraction rates of variational posteriors under certain assumptions appear in several recent works Zhang and Gao (2020); Alquier and Ridgway (2020); Bhattacharya et al. (2023); Yang et al. (2020) and contraction rates as well as model selection consistency in mixtures are considered in Chérief-Abdellatif (2020).
Several works focus on properties of VI approximation in sparse models: oracle contraction rates of a mean-field VI in the case of high-dimensional regression are established in Ray and Szabó (2022) and in the case of the neural networks with spike-and-slab priors, contraction rates of variational posteriors are studied in Bai et al. (2020). More recently, near-optimal contraction rates of the mean-field VI approximations are obtained in the context of neural networks with heavy-tailed priors on the weights Castillo and Egels (2024). We refer readers interested in the caveats of VI to Yao et al. (2018).
@@
A common choice for $\mathcal{F}$ is the mean-field family on a partition $\{\bm{\theta}_{1},\ldots,\bm{\theta}_{J}\}$ of $\bm{\theta}$, assuming that the variational posterior factorizes over (blocks) of latent variables: $q(\bm{\theta})=\prod_{j=1}^{J}q_{j}(\bm{\theta}_{j}),$ where $J\leq\text{dim}(\bm{\theta})$. Without any further parametric assumptions, it has been shown Hinton and van Camp (1993); MacKay (1995); Jordan et al. (1999) that the optimal choice for each product component $q_{j}$ is
@@
where the above expectation is taken with respect to $\prod_{j^{\prime}\neq j}q_{j^{\prime}}(\bm{\theta}_{j^{\prime}})$. The process of sweeping through the components of the partition and updating one at a time via Equation 11 is known as coordinate ascent variational inference (CAVI). Wand et al. (2011) studies CAVI’s performance and improvement techniques in the case of elaborate distributions. Limitations of the mean-field assumption can be found in Wand et al. (2011); Neville et al. (2014); Coker et al. (2022).
@@
$q(\mathbf{a},\bm{\gamma},\bm{\omega},\mathbf{W},\mathbf{b},\bm{\eta},\bm{\psi},\bm{\tau})=q(\mathbf{a})q(\bm{\gamma})q(\bm{\omega})q(\mathbf{W},\mathbf{b})q(\bm{\eta})q(\bm{\psi})q(\bm{\tau}).$
Note that the assumption on the family above could be referred to as the strucutred mean-field assumtion. Importantly, unlike existing variational algorithms for BNNs, we do not make any assumptions on independence of parameters between layers. The calculation of each component of the variational posterior is given in the Appendix A, where using Equation 11 we obtain the following update steps.
@@
$q(\bm{\tau})$ $=\prod_{l}^{L+1}\text{GIG}\left(\tau_{l}\;\mid\;\hat{\nu}_{\text{glob},l},\hat{\delta}_{\text{glob},l},\lambda_{\text{glob}}\right),$
@@
$q(\bm{\psi})$ $=\prod_{l}^{L+1}\prod_{d}^{D_{l}}\prod_{d^{\prime}}^{D_{l-1}}\text{GIG}\left(\psi_{l,d,d^{\prime}}\;\mid\;\hat{\nu}_{\text{loc},l,d,d^{\prime}},\hat{\delta}_{\text{loc},l,d,d^{\prime}},\lambda_{\text{loc},l}\right),$
@@
$q(\bm{\eta})=\prod_{l}^{L+1}\prod_{d}^{D_{l}}\text{IG}(\eta_{l,d}^{2}\mid\alpha_{l,d},\beta_{l,d}),$
@@
$\beta_{l,d}=\beta^{h}_{0}+\frac{1}{2}\sum_{n}^{N}\left(\mathbb{E}\left[a_{n,l,d}\right]-\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\right]\right)^{2}+\mathbb{E}\left[a^{2}_{n,l,d}\right]-\mathbb{E}\left[a_{n,l,d}\right]^{2}$
@@
the weights and biases are independent across layers and within layer, independent across the $D_{l}$ regression problems, with a Gaussian variational posterior:
$q(\mathbf{b},\mathbf{W})=\prod_{l}^{L+1}\prod_{d}^{D_{l}}\text{N}\left(\widetilde{\mathbf{W}}_{l,d}\mid\mathbf{m}_{l,d},\mathbf{B}_{l,d}\right),$
@@
$\mathbf{m}_{l,d}^{T}$ $=\mathbf{B}_{l,d}\left(\sum_{n}^{N}\mathbb{E}\left[(\eta_{l,d})^{-2}\right]\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[a_{n,l,d}\widetilde{\mathbf{a}}_{n,l-1}\right]+\frac{1}{T}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\right]\left(\mathbb{E}\left[\gamma_{n,l,d}\right]-\frac{1}{2}\right)\right),$
@@
$\mathbf{B}^{-1}_{L+1,d}$ $=\mathbf{D}^{-1}_{L+1,d}+\mathbb{E}\left[(\eta_{L+1,d})^{-2}\right]\sum_{n}^{N}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,L+1}\widetilde{\mathbf{a}}_{n,L+1}^{T}\right],$
$\mathbf{m}_{L+1,d}^{T}$ $=\mathbf{B}_{L+1,d}\mathbb{E}\left[(\eta_{L+1,d})^{-2}\right]\left(\sum_{n}^{N}y_{n}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,L+1}\right]\right),$
@@
$\rho_{n,l,d}=$ $\sigma\left(-\frac{\mathbb{E}\left[\eta_{l,d}^{-2}\right]}{2}\operatorname{Tr}\left(\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}^{T}\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\widetilde{\mathbf{a}}_{n,l-1}^{T}\right]\right)+\mathbb{E}\left[\eta_{l,d}^{-2}\right]\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}a_{n,l,d}\right]+\frac{1}{T}\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\right]\right).$
We illustrate the variational posterior of $\bm{\gamma}$ for the toy example of Section 4.1 in Figure 4.
@@
Figure 4 illustrates on the toy example of Section 4.1 how the variational posterior of the stochastic activations (middle) resembles a smoothed, noisy ReLU. Due the independence assumption between the stochastic and binary activations, the potentially bimodal bow tie distribution (Equation 2) is approximated with a unimodal Gaussian in the variational framework, which may better approximate the true posterior when the temperature is not too large relative to the noise (see Figure 1). In addition, the proposed approximation has the advantage of avoiding explicit assumptions of independence between layers, allowing to capture the dependence between the stochastic activations of across layers visualized for the toy example in Figure 4 (right).
Figure 4: Variational posterior for $\rho_{n,1,d}$ (on the left) and $a_{n,1,d}$ (in the middle), joint distribution of $(a_{n,1,d},a_{n,2,d^{\prime}})$ (on the right) in the case of the toy example of Section 4 and particular values of $d,d^{\prime}$.
The corresponding optimization objective, i.e. the ELBO in Equation 10, is available in the closed form and provided in the Appendix B.
@@ 3.2 VI with EM
The hyperparameters can play a crucial role in Bayesian neural networks. When dealing with the sparsity-inducing priors setting an excessively large scale parameter weakens the shrinkage effects, whilst choosing a scale parameter that is too small may wipe out the effects of the important hidden nodes. Manually picking suitable values is challenging, and instead, we seek a more efficient strategy, utilizing the similarity between the variational and expectation-maximization (EM) algorithms. Specifically, we investigate the hybrid scheme combining VI with an EM step Osborne et al. (2022) so that the steps of the CAVI algorithm proceed with the EM update to set the hyperparameter for global shrinkage variable $\bm{\tau}$. Due to weak identifiability, we do not jointly update global and local hyperparameters. Let $h_{\text{glob}}$ represent $\delta_{\text{glob}}$ or $\lambda_{\text{glob}}$ and consider the ELBO treated as a function of $h_{\text{glob}}$, then the optimal values as approximate MAP estimates are:
where
In the case of the IG priors, one’s aim is to set optimal $\delta_{\text{glob}}$, in the case of the Gamma and IGauss priors, the parameters of interest are $\lambda_{\text{glob}}$. We provide specific examples of the shrinkage parameters and the corresponding optimal values in Section D.2. The result of combining CAVI and the EM algorithm is described in Algorithm 1.

### v3

@@ 2.3 Polya-Gamma Data Augmentation
As in (Smith et al., 2021), we employ Polya-Gamma data augmentation (Polson et al., 2013) to render the model conditionally linear and Gaussian. First, recall the definition of the Polya-Gamma distribution with parameters $b>0$ and $c\in\mathbb{R}$, denoted $\text{PG}(b,c)$. The random variable $X\sim\text{PG}(b,c)$ if
@@
where $\kappa=a-b/2$ and $p(\omega)=PG(\omega|b,0)$. The integral is a Gaussian kernel, thus if $z=\mathbf{w}^{T}\mathbf{x}$, conditioned on the latent variable $\omega$, $\mathbf{w}$ has a Gaussian distribution and conditioned on $\mathbf{w}$, $\omega$ has a PG distribution. While to sample from the PG distribution, one can use the alternating series method of (Devroye, 2006), all finite moments of the PG random variables are available in closed form, and that becomes useful for expectation-maximization or variational Bayes algorithms. Specifically, for $c>0$
@@ 2.4 Augmented Model
The model in Section 2.1 augmented with stochastic activations $\mathbf{a}=(\mathbf{a}_{n,l})$ and binary activations $\bm{\gamma}=(\bm{\gamma}_{n,l})$ is:
@@
The covariance matrices are assumed to be diagonal $\bm{\Sigma}_{l}=\text{diag}(\eta_{l,1}^{2},\ldots\eta_{l,D_{l}}^{2})$, with variances denoted by $\bm{\eta}_{l}=(\eta_{l,1}^{2},\ldots\eta_{l,D_{l}}^{2})$. Additionally, we assume conjugate priors for the variances $\eta_{l,d}^{2}\overset{iid}{\sim}\text{IG}(\alpha^{h}_{0},\beta^{h}_{0})$ for $l=1,\ldots,L$ and $\eta_{L+1,d}^{2}\overset{iid}{\sim}\text{IG}(\alpha_{0},\beta_{0})$ and for the biases $b_{l,d}\overset{iid}{\sim}\text{N}(0,s^{2}_{0})$. Here, we consider different prior parameters $\alpha_{0}^{h},\beta_{0}^{h}$ for the variance terms associated to the hidden layers in comparison to the parameters $\alpha_{0},\beta_{0}$ for the final layer. In particular, $\alpha_{0},\beta_{0}$ are chosen to reflect prior knowledge in the noise, while $\alpha_{0}^{h},\beta_{0}^{h}$ are chosen so that the prior concentrates on small values and realizations of the stochastic activation function are more similar to the ReLU.
@@
×$\times\prod_{d=1}^{D_{l}}\exp\left(\frac{\kappa_{n,l,d}z_{n,l,d}}{T}\right)\exp\left(-\frac{\omega_{n,l,d}z_{n,l,d}^{2}}{2T^{2}}\right)p(\omega_{n,l,d})$
×$\times\prod_{d=1}^{D_{L+1}}\text{IG}(\eta^{2}_{L+1,d}\mid\alpha_{0},\beta_{0})\times\prod_{l=1}^{L}\prod_{d=1}^{D_{l}}\text{IG}(\eta^{2}_{l,d}\mid\alpha_{0}^{h},\beta_{0}^{h})\times\prod_{n=1}^{N}\prod_{l=1}^{L}\prod_{d=1}^{D_{l}}\text{Bern}\left(\gamma_{n,l,d}\mid\sigma\left(\frac{z_{n,l,d}}{T}\right)\right)$
@@ 3.1 Variational Bayes
While Markov chain Monte Carlo is considered the gold-standard tool for approximating posterior distributions in Bayesian modeling due to its asymptotic guarantees, MCMC algorithms can be prohibitively slow when the model dimension and sample size are large. Instead, we focus on variational inference, an alternative fast, approximate Bayesian inference method that has gained popularity in the literature (Ormerod and Wand, 2010; Zhang et al., 2018), due to both the explosion in the amount of data collected and the use of highly parametrized models for increased flexibility. VI has been shown to yield reasonably accurate approximations in several problems as well as desirable frequentist properties. Namely, consistency and asymptotic normality of the variational posterior expectation are established in (Wang and Blei, 2019), and theoretical guarantees for optimal contraction rates of variational posteriors under certain assumptions appear in several recent works (Zhang and Gao, 2020; Alquier and Ridgway, 2020; Bhattacharya et al., 2023; Yang et al., 2020; Ray and Szabó, 2022). We refer readers interested in the caveats of VI to (Yao et al., 2018; Barber et al., 2011; Hron et al., 2018).
@@
A common choice for $\mathcal{F}$ is the nonparametric, mean-field family on a partition $\{\bm{\theta}_{1},\ldots,\bm{\theta}_{J}\}$ of $\bm{\theta}$, assuming that the variational posterior factorizes over (blocks) of latent variables: $q(\bm{\theta})=\prod_{j=1}^{J}q_{j}(\bm{\theta}_{j}),$ where $J\leq\text{dim}(\bm{\theta})$. Without any further parametric assumptions, it has been shown (Hinton and van Camp, 1993; MacKay, 1995; Jordan et al., 1999) that the optimal choice for each product component $q_{j}$ is
@@
where the above expectation is taken with respect to $\prod_{j^{\prime}\neq j}q_{j^{\prime}}(\bm{\theta}_{j^{\prime}})$. The process of sweeping through the components of the partition and updating one at a time via Equation 11 is known as coordinate ascent variational inference (CAVI). (Wand et al., 2011) studies CAVI’s performance and improvement techniques in the case of elaborate distributions. Limitations of the mean-field assumption can be found in (Wand et al., 2011; Neville et al., 2014; Coker et al., 2022).
@@
$q(\mathbf{a},\bm{\gamma},\bm{\omega},\mathbf{W},\mathbf{b},\bm{\eta},\bm{\psi},\bm{\tau})=q(\mathbf{a})q(\bm{\gamma})q(\bm{\omega})q(\mathbf{W},\mathbf{b})q(\bm{\eta})q(\bm{\psi})q(\bm{\tau}).$ (12)
Note that the assumption on the family above could be referred to as the structured mean-field assumption. Importantly, unlike existing variational algorithms for BNNs, we do not make any assumptions on independence of parameters between layers. The calculation of each component of the variational posterior is given in the Appendix A, where using Equation 11 we obtain the following update steps.
@@
$q(\bm{\tau})$ $=\prod_{l}^{L+1}\text{GIG}\left(\tau_{l}\;\mid\;\hat{\nu}_{\text{glob},l},\hat{\delta}_{\text{glob},l},\lambda_{\text{glob}}\right),$ (13)
@@
$q(\bm{\psi})$ $=\prod_{l}^{L+1}\prod_{d}^{D_{l}}\prod_{d^{\prime}}^{D_{l-1}}\text{GIG}\left(\psi_{l,d,d^{\prime}}\;\mid\;\hat{\nu}_{\text{loc},l,d,d^{\prime}},\hat{\delta}_{\text{loc},l,d,d^{\prime}},\lambda_{\text{loc},l}\right),$ (14)
@@
$q(\bm{\eta})=\prod_{l}^{L+1}\prod_{d}^{D_{l}}\text{IG}(\eta_{l,d}^{2}\mid\alpha_{l,d},\beta_{l,d}),$ (15)
@@
$\beta_{l,d}$ $=\beta^{h}_{0}+\frac{1}{2}\sum_{n}^{N}\left(\mathbb{E}\left[{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\right]-\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\right]\right)^{2}+\mathbb{E}\left[a^{2}_{n,l,d}\right]-\mathbb{E}\left[{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\right]^{2}$
@@
the weights and biases are independent across layers and within layers, independent across the $D_{l}$ regression problems, with a Gaussian variational posterior:
$q(\mathbf{b},\mathbf{W})=\prod_{l}^{L+1}\prod_{d}^{D_{l}}\text{N}\left(\widetilde{\mathbf{W}}_{l,d}\mid\mathbf{m}_{l,d},\mathbf{B}_{l,d}\right),$ (16)
@@
$\mathbf{m}_{l,d}^{T}$ $=\mathbf{B}_{l,d}\left(\sum_{n}^{N}\mathbb{E}\left[(\eta_{l,d})^{-2}\right]\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\widetilde{\mathbf{a}}_{n,l-1}\right]+\frac{1}{T}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\right]\left(\mathbb{E}\left[\gamma_{n,l,d}\right]-\frac{1}{2}\right)\right),$
@@
[…]
$\mathbf{m}_{L+1,d}^{T}$ =$𝐿\displaystyle=\mathbf{B}_{L+1,d}\mathbb{E}\left[(\eta_{L+1,d})^{-2}\right]\left(\sum_{n}^{N}y_{n}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,{\color[rgb]{0,0,0}{L}}}\right]\right),$
@@
$\rho_{n,l,d}=$ $\sigma\left(-\frac{\mathbb{E}\left[\eta_{l,d}^{-2}\right]}{2}\operatorname{Tr}\left(\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}^{T}\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\widetilde{\mathbf{a}}_{n,l-1}^{T}\right]\right)+\mathbb{E}\left[\eta_{l,d}^{-2}\right]\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\right]+\frac{1}{T}\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\right]\right).$
The parameters $\rho_{n,l,d}$ represent the posterior probability that the node is active and are illustrated for the toy example of Section 4.1 in Figure 4.
@@
Figure 4 illustrates on the toy example of Section 4.1 how the variational posterior of the stochastic activations (middle) resembles a smoothed, noisy ReLU. Due the independence assumption between the stochastic and binary activations, the potentially bimodal bow tie distribution (Equation 2) is approximated with a unimodal Gaussian in the variational framework, which may better approximate the true posterior when the temperature is not too large relative to the noise (see Figure 1). In addition, the proposed approximation has the advantage of avoiding explicit assumptions of independence between layers, enabling it to capture dependencies between stochastic activations across layers, as illustrated for the toy example in Figure 4 (right).
Figure 4: An illustration of the variational posterior of the binary and stochastic activations. The variational posterior of $\gamma_{n,1,d}$ (on the left) and $a_{n,1,d}$ (in the middle), both as a function of $x_{n,1}$ across all observations, along with the joint distribution of $(a_{n,1,d},a_{n,2,d^{\prime}})$ (on the right) in the case of the toy example of Section 4 for particular values of $d,d^{\prime},n$.
The corresponding optimization objective, i.e. the ELBO in Equation 10, is available in closed form and provided in the Appendix B.
@@ 3.2 VI with EM
The hyperparameters can play a crucial role in Bayesian neural networks. When dealing with the sparsity-inducing priors setting an excessively large scale parameter weakens the shrinkage effects, whilst choosing a scale parameter that is too small may wipe out the effects of the important hidden nodes. Manually picking suitable values is challenging, and instead, we seek a more efficient strategy, utilizing the similarity between the variational and expectation-maximization (EM) algorithms. Specifically, we investigate the hybrid scheme combining VI with an EM step (Osborne et al., 2022) so that the steps of the CAVI algorithm proceed with the EM update to set the hyperparameter for global shrinkage variable $\bm{\tau}$. Due to weak identifiability, we do not jointly update global and local hyperparameters. Let $h_{\text{glob}}$ represent $\delta_{\text{glob}}$ or $\lambda_{\text{glob}}$ and consider the ELBO treated as a function of $h_{\text{glob}}$, then the optimal values as approximate MAP estimates are:
where
In the case of the IG priors, one’s aim is to set optimal $\delta_{\text{glob}}$, in the case of the Gamma and IGauss priors, the parameters of interest are $\lambda_{\text{glob}}$. We provide specific examples of the shrinkage parameters and the corresponding optimal values in Section E.2.
The resulting algorithm (described in Algorithm 1) scales linearly with the number of hidden layers and the number of samples but cubically with the number of hidden units; the computational complexities corresponding to individual factors of the variational family are provided in Appendix A. In the following subsections, we discuss two strategies to improve scalability. First, to handle large $N$, CAVI can be combined with subsampling through stochastic VI (Hoffman et al., 2013), and second, a post-processing node selection algorithm is proposed to obtain a sparse variational posterior for faster predictive inference.
3.3 Stochastic Variational Inference
At each iteration, CAVI has to cycle through the entire data set, which can be computationally expensive and inefficient for large sample sizes. An alternative to coordinate ascent is gradient-based optimization, which extends the algorithm by employing stochastic variational inference (SVI Hoffman et al., 2013) and subsampling. Forst. recall that the variational posterior of the latent variables $\mathbf{a},\bm{\gamma},\bm{\omega}$ factorizes across data points, that is
$q(\mathbf{a})q(\bm{\gamma})q(\bm{\omega})=\prod_{n=1}^{N}q(\mathbf{a}_{n})q(\bm{\gamma}_{n})q(\bm{\omega}_{n}).$
To highlight the need for stochastic VI, we observe that for each layer $l=1,\ldots,L$ the ordinary coordinate ascent needs to iterate through $N$ local variational parameters corresponding to variables $\mathbf{a}_{n,l},\bm{\gamma}_{n,l}$ and $\bm{\omega}_{n,l}$ with $\bm{\gamma}_{n,l}$ and $\bm{\omega}_{n,l}$ both having $D_{l}$ variational parameters to update, and $\mathbf{a}_{n,l}$ having of order $D_{l}^{2}$ variational parameters. For large $N$, this can be computationally burdensome. To overcome this, stochastic VI uses the coordinate ascent variational updates obtained in Section 3.1 to set the local parameters for a subset of points, and the intermediate global parameters are obtained by following the noisy natural gradient of ELBO (Hoffman et al., 2013).
At each iteration $t$, the SVI algorithm first uniformly samples a collection of indices $S(t)$ without replacement, where the computational gains are obtained as long as $|S(t)|\ll N$. For each index $s\in S(t)$, the local variational parameters are optimized in a coordinate ascent manner with updates in Equations 19, 17 and 18, and the convergence is monitored with noisy estimates of the local ELBO, that is
$\mathbb{E}\left[\log p(\mathbf{y},\mathbf{a},\bm{\gamma},\bm{\omega}|\mathbf{W},\mathbf{b},\bm{\Sigma})\right]-\mathbb{E}\left[\log q(\mathbf{a})\right]-\mathbb{E}\left[\log q(\bm{\gamma})\right]-\mathbb{E}\left[\log q(\bm{\omega})\right],$
the estimate of which is provided Appendix C. Next, the global parameters, $\mathbf{W},\mathbf{b}$ and $\bm{\eta}$, are updated via a linear combination of previous values and intermediate updates
$(1-\ell_{t})\text{parameter}^{(t-1)}+\ell_{t}\times\text{intermediate parameter}^{(t)},$ (20)
where $\ell_{t}$ is the learning rate and the intermediate parameters are obtained via Equations 15 and 16 but with the sufficient statistics, which involve sums over $N$, replaced with the scaled sums over $S(t)$. Additional details and the step-by-step algorithm are provided in Appendix C. The convergence of the stochastic gradient descent depends on the choice of the step sizes $\ell_{t}$ in the Robbins-Monro sequence (Robbins and Monro, 1951), and we follow (Hoffman et al., 2013), who set
$\ell_{t}=(1+t)^{-k},\;k\in(0.5,1],$ (21)
where $k$ is the forgetting rate. We monitor the convergence of the algorithm by obtaining noisy estimates of the ELBO (provided in Appendix C) where the sums over $n=1,\ldots,N$ are replaced with the scaled sums over $n\in S$.